OpenStudy (anonymous):

Find the slope of the line passing through E(2,5) and F(7,9).

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (anonymous):

just put the values

OpenStudy (anonymous):

oh so its m=(9-5)+(7-2)

OpenStudy (anonymous):

9-5/7-2 4/5=slope

OpenStudy (anonymous):

so its m=(9-5)/(7-2)

OpenStudy (anonymous):

That's right @laxplayer

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@J.L.

OpenStudy (anonymous):

@nishika17

OpenStudy (anonymous):

i have another problem

OpenStudy (anonymous):

go ahead

OpenStudy (anonymous):

Find the slope of the line passing through Q(5,-1) and P (5,5).

OpenStudy (anonymous):

so do i do the same thing as before

OpenStudy (anonymous):

its the same as before

OpenStudy (anonymous):

Yup! Use the same formula

OpenStudy (anonymous):

oh okay thats what i thought

OpenStudy (anonymous):

@laxplayer Just curious...you're good at geometry, but you're asking an algebra problem? Just wondering... \[m = \frac{y_1 - y_2}{x_1 - x_2}\]

OpenStudy (anonymous):

Your slope should be 6

OpenStudy (anonymous):

well iam good at some stuff ad bad at other stuff @Calcmathlete

OpenStudy (anonymous):

Any other questions?

OpenStudy (anonymous):

these are the answers

OpenStudy (anonymous):

1 undefined 0

OpenStudy (anonymous):

no 6???

OpenStudy (anonymous):

slope is undefined

OpenStudy (anonymous):

how

OpenStudy (anonymous):

oh i cant divide by zero duh

OpenStudy (anonymous):

y2-y1/x2-x1 (5--1)/(5-5) (5+1)/(0) Slope=6/0-undefined

OpenStudy (anonymous):

Yup...denominator can't equal 0.

OpenStudy (anonymous):

yup basic math i forgot

OpenStudy (anonymous):

What is the slope of the side AB of the triangle ΔABC with vertices A (-3,3), B(2,4) and C(1,-2)?

OpenStudy (anonymous):

Then how do i do this the same way

OpenStudy (anonymous):

Find the slope of AB

OpenStudy (anonymous):

(4-3)/(2+3)

OpenStudy (anonymous):

oh so it is the same way

OpenStudy (anonymous):

1/5 or .2

OpenStudy (anonymous):

Thanks i got no more question

OpenStudy (anonymous):

Thanks alot @J.L. and @Calcmathlete